OpenStudy (anonymous):

Find the particular solution for 2x (dy/dx)-ln x^2 that satisfies the initial condition y(1) = 2. Explain your steps.

OpenStudy (anonymous):

\[2x \frac{ dy }{ dx }-\ln x ^{2}\] equation looks like that

ganeshie8 (ganeshie8):

separation of variables

OpenStudy (anonymous):

ok i need help with the whole problem

ganeshie8 (ganeshie8):

sorry was afk...

ganeshie8 (ganeshie8):

\(\large 2x \frac{ dy }{ dx }-\ln x^{2}\) \(\large 2x \frac{ dy }{ dx } = \ln x^{2}\) \(\large 2 ~dy = \frac{\ln x^{2}}{x}dx \)

ganeshie8 (ganeshie8):

Integrate both sides

ganeshie8 (ganeshie8):

\(\large \int 2 ~dy = \int \frac{\ln x^{2}}{x}dx \)

ganeshie8 (ganeshie8):

evaluate

OpenStudy (anonymous):

ok thank you!

ganeshie8 (ganeshie8):

u sure.. u can manage the rest ha ? :)

ganeshie8 (ganeshie8):

integrating right side can be tricky... u need to use "u sub"

OpenStudy (anonymous):

no i cant hahah

ganeshie8 (ganeshie8):

\(\large \int 2 ~dy = \int \frac{\ln x^{2}}{x}dx \) \(\large 2\int ~dy = \int \frac{2\ln x}{x}dx \) \(\large 2y = 2\int \frac{\ln x}{x}dx \)

OpenStudy (anonymous):

is that it?

ganeshie8 (ganeshie8):

substitute \(\ln x = t\) \(\frac{1}{x} dx = dt\) right side beocmes : \(\large 2\int t~ dt \) \(\large 2 \frac{t^2}{2} + C \) \(\large t^2+ C \) \(\large (\ln x)^2+ C \)

ganeshie8 (ganeshie8):

plug that in right hand side

ganeshie8 (ganeshie8):

\(\large 2y = (\ln x)^2 + C\)

ganeshie8 (ganeshie8):

you need to plugin the initial condition (1, 2) and solve \(C\) : \(\large 2*2 = (\ln 1)^2 + C \implies C = 4\)

ganeshie8 (ganeshie8):

So, the solution is : \(\large 2y = (\ln x)^2 + 4\)

OpenStudy (anonymous):

thank you so much for your help! This'll definitely help me with similar problems

ganeshie8 (ganeshie8):

np :) good luck !

ganeshie8 (ganeshie8):

i made a mistake earlier

ganeshie8 (ganeshie8):

nvm, it is correct :)